OpenStudy (anonymous):

Solve without using a calculator step by step plz sin 30 degree csc 30 degree

OpenStudy (anonymous):

@Hero

OpenStudy (raden):

hint : csc30 = 1/sin30

OpenStudy (anonymous):

to be honest i know that csc =1/sin and sin= 1/csc and now im stock

OpenStudy (raden):

so, how do u simplify of sin30 * 1/sin30 is .....

OpenStudy (raden):

it is like u multiply 2 * 1/2, 3 * 1/3, 4 * 1/4, etc...

OpenStudy (anonymous):

oooh

OpenStudy (raden):

so, = ?

Directrix (directrix):

Do some reading about reciprocal trig functions at this site: http://www.regentsprep.org/Regents/math/algtrig/ATT1/trigsix.htm

OpenStudy (anonymous):

use the unit circle. sin(30)= 1/2

Directrix (directrix):

Is this a new problem? --> use the unit circle. sin(30)= 1/2

OpenStudy (anonymous):

this is so confusing lol

OpenStudy (anonymous):

it is like this \[\sin(x)\times \csc(x)=\frac{\sin(x)}{\sin(x)}=1\]

OpenStudy (anonymous):

the 30 degrees part is irrelevant it is one no matter what

OpenStudy (anonymous):

well can you please solve this in a sequential order so i can understand it

Directrix (directrix):

Have you studied any trigonometric identities yet? Have you learned that sine of 30 degrees is 1/2 yet? Have you learned the 30-60-90 theorem yet? I ask because we need to know where you are in your study of trig.

OpenStudy (anonymous):

i did first one

OpenStudy (anonymous):

im now on inverse

Directrix (directrix):

I think maybe that drawing is what you want although I am not sure.

OpenStudy (anonymous):

at the risk of repeating myself, the 30 degrees in the problem is irrelevant

Directrix (directrix):

@magepker728 wanted a unit circle demonstration.

OpenStudy (anonymous):

i see with problems like this i need to think of the unit circle...